# AI-Augmented Thyroid Scintigraphy for Robust Classification

### v1

@@ 4.1 Training set-up
Conventional augmentation: We applied a randomized transformation pipeline that included rotation (±15°), horizontal flipping, translation (±10%), scaling (0.8–1.2×), and Gaussian noise addition ($\sigma$ = 0.001–0.01).
Stable diffusion augmentation: We used a fine-tuning setup with mixed precision (fp16), a resolution of 128×128, exponential moving average, gradient accumulation with four steps, and a batch size of 1, optimizing for 50,000 steps at a learning rate of 1e-5. During inference, we evaluated three approaches: 1) image and prompt to image (SD1), 2) prompt to image (SD2), and 3) mask and prompt to image (SD3).
Flow matching augmentation: For FM, the model was optimized using the Adam optimizer with a learning rate of 1e-4 for 200 epochs. Our approach leverages FM with optimal transport to align predicted data flows with the target distribution. Class conditioning is achieved by incorporating a one-hot encoded vector via cross-attention, while mask conditioning is implemented through a parallel control network integrated via residual connections. During inference, we assessed guided generation using a combination of mask and class conditioning.
This resulted in 15 distinct training strategies for the ResNet18 classifier: 1) O, 2) CA, 3) SD1, 4) SD2, 5) SD3, 6) FM, 7) O+CA, 8) O+SD1, 9) O+SD2, 10) O+SD3, 11) O+FM, 12) O+CA+SD1, 13) O+CA+SD2, 14) O+CA+SD3, and 15) O+CA+FM.
ResNet18: We trained a ResNet18 classifier, initializing it with ImageNet1K pre-trained weights. The first convolutional layer was modified to retain a 3×3 kernel, and the fully connected layer was replaced with a dropout layer (0.2) followed by a linear layer matching the number of classes. The model was trained for 300 epochs using the Adam optimizer (learning rate = 1e-4, weight decay = 1e-5) with a cross-entropy loss function. A learning rate scheduler (ReduceLROnPlateau) was applied, reducing the rate by a factor of 0.8 if validation loss plateaued for 10 epochs, with a minimum learning rate of 2e-5. Furthermore, the training and validation split was set at a 9:1 ratio with stratification.
@@ 4.2 Evaluation metrics
Augmentation metrics: Fréchet Inception Distance (FID) [10] and Kernel Inception Distance (KID) [5] metrics have been used for class-wise and overall comparisons between generated and original images. In both cases, 200 images per class were selected from each dataset for evaluation.
@@
Classification metrics: The classification performance was evaluated on an external dataset using metrics, including precision, recall, F1-score, accuracy, and the area under the Receiver Operating Characteristic curve (ROC AUC). Given that it was a multiclass task, we applied various averaging techniques encompassing micro, macro, and weighted to provide a thorough evaluation across all classes.
@@
Statistical method: We compared different strategies using bootstrapping with 1,000 repetitions and sampling with replacement. Accuracy distributions were analyzed, and pairwise comparisons were made using the Wilcoxon rank sum test, considering p-values < 0.5 as statistically significant.
@@
GradCam: We used Gradient-weighted Class Activation Mapping (GradCAM) on the trained ResNet18 model to generate heatmaps, visualizing the image regions that influenced the classifier’s decisions.
5 Results and analysis
Figure 1 presents one sample per class from the original dataset alongside examples generated by different augmentation methods used in this study. Additionally, it includes samples from the external dataset with their corresponding Grad-CAM visualizations using the O+FM+CA model, highlighting the model’s focus during prediction. The analysis of FID and KID metrics (Table 3) reveals significant performance differences. SD3 performs the worst with the highest FID (17.99) and KID Mean (33.59), indicating poor image quality and large distribution shifts. SD1 and SD2 perform better, with SD2 achieving the lowest FID (3.88), though its TI class shows a high KID Mean (30.02). FM outperforms all, with the best FID (0.66) and minimal KID Mean, indicating its generated images closely match real data. Table 4 and Table 5 highlight the impact of different model configurations on classification performance. The inclusion of FM and CA components consistently improves performance across all metrics, with O+FM+CA and O+FM achieving the highest F1-scores. Notably, O+FM provides the best balance between precision and recall, while FM alone offers strong precision but slightly lower recall. In terms of accuracy and AUC, O+FM outperforms other models, reaching an accuracy of 0.78 and the highest AUC values. SD3, which exhibits low recall and AUC, lags behind other configurations. Additionally, the Wilcoxon rank sum test (Figure 2) confirms that O+FM and O+FM+CA consistently outperform other models, while SD1, which incorporates both prompt and image during inference, outperforms SD2 and SD3. SD2, which lacks image guidance, and SD3, which struggles with masked inputs, show suboptimal results.
Table 3: Comparison of FID and KID Mean (Std < 1e-5 for all) for SD and FM Methods.
Class SD1 SD2 SD3 FM
FID $\downarrow$ KID $\downarrow$ FID $\downarrow$ KID $\downarrow$ FID $\downarrow$ KID $\downarrow$ FID $\downarrow$ KID $\downarrow$
DG 4.75 6.77 2.94 2.31 10.30 15.40 0.96 0.80
NG 6.22 10.28 3.27 2.66 22.07 38.34 0.74 0.95
NL 4.56 7.14 1.75 2.05 18.68 36.74 0.85 2.08
TI 2.73 3.01 9.24 30.02 22.37 50.50 1.97 3.39
Overall 4.17 4.99 3.88 2.61 17.99 33.59 0.66 0.83
@@ 5 Results and analysis
Figure 1: Examples of original and augmented images for each class using different methods. Grad-CAM visualizations from the O+FM+CA model are also shown for external dataset samples, highlighting the model’s focus during prediction.
Table 4: Class-wise performance metrics of the classification. (Support: DG: 130, NG: 95, NL:74, and TI: 100)
Method Precision Recall F1-score
(DG/NG/NL/TI) (DG/NG/NL/TI) (DG/NG/NL/TI)
O 0.70/0.65/0.53/0.96 0.72/0.86/0.31/0.91 0.71/0.74/0.39/0.93
CA 0.75/0.70/0.55/0.92 0.62/0.77/0.62/0.97 0.68/0.73/0.58/0.95
SD1 0.80/0.74/0.53/0.92 0.72/0.71/0.59/0.98 0.76/0.72/0.56/0.95
SD2 0.76/0.57/0.45/0.95 0.70/0.68/0.53/0.74 0.73/0.62/0.49/0.83
SD3 0.88/0.66/0.42/0.58 0.44/0.35/0.64/1.00 0.58/0.46/0.50/0.74
FM 0.70/0.54/0.48/0.96 0.67/0.64/0.54/0.76 0.69/0.59/0.51/0.85
O+CA 0.73/0.59/0.61/0.99 0.72/0.83/0.41/0.89 0.72/0.69/0.49/0.94
O+SD1 0.74/0.67/0.55/0.96 0.70/0.85/0.46/0.89 0.72/0.75/0.50/0.92
O+SD2 0.78/0.60/0.61/0.98 0.67/0.87/0.49/0.89 0.72/0.71/0.54/0.93
O+SD3 0.76/0.62/0.59/0.95 0.66/0.87/0.47/0.89 0.71/0.73/0.53/0.92
O+FM 0.74/0.74/0.68/0.93 0.81/0.76/0.54/0.95 0.77/0.75/0.60/0.94
O+SD1+CA 0.75/0.64/0.54/0.96 0.66/0.82/0.50/0.91 0.70/0.72/0.52/0.93
O+SD2+CA 0.74/0.65/0.62/0.94 0.66/0.82/0.54/0.91 0.70/0.73/0.58/0.92
O+SD3+CA 0.72/0.63/0.58/0.95 0.63/0.82/0.53/0.90 0.67/0.72/0.55/0.92
O+FM+CA 0.79/0.68/0.63/0.97 0.73/0.84/0.58/0.90 0.76/0.75/0.61/0.93
Table 5: Averaging performance metrics of the classification. (Support: DG+NG+NL+TI=399)
Method AUC Precision Recall F1-score
(mic/mac/wei) (mic/mac/wei) (mic/mac/wei) (mic/mac/wei)
O 0.92/0.91/0.91 0.73/0.71/0.72 0.73/0.70/0.73 0.73/0.69/0.71
CA 0.93/0.92/0.92 0.74/0.73/0.75 0.74/0.74/0.74 0.74/0.74/0.74
SD1 0.93/0.92/0.92 0.76/0.74/0.76 0.76/0.75/0.76 0.76/0.75/0.76
SD2 0.88/0.87/0.88 0.67/0.68/0.70 0.67/0.66/0.67 0.67/0.67/0.68
SD3 0.82/0.86/0.86 0.59/0.63/0.67 0.59/0.61/0.59 0.59/0.57/0.58
FM 0.89/0.88/0.88 0.66/0.67/0.69 0.66/0.65/0.66 0.66/0.66/0.67
O+CA 0.92/0.92/0.92 0.73/0.73/0.74 0.73/0.71/0.73 0.73/0.71/0.73
O+SD1 0.92/0.92/0.92 0.74/0.73/0.74 0.74/0.73/0.74 0.74/0.72/0.74
O+SD2 0.92/0.92/0.92 0.74/0.74/0.76 0.74/0.73/0.74 0.74/0.73/0.74
O+SD3 0.92/0.92/0.92 0.73/0.73/0.74 0.73/0.72/0.73 0.73/0.72/0.73
O+FM 0.95/0.93/0.94 0.78/0.77/0.78 0.78/0.76/0.78 0.78/0.77/0.78
O+SD1+CA 0.93/0.92/0.93 0.73/0.72/0.74 0.73/0.72/0.73 0.73/0.72/0.73
O+SD2+CA 0.92/0.92/0.92 0.74/0.73/0.74 0.74/0.73/0.74 0.74/0.73/0.74
O+SD3+CA 0.92/0.91/0.92 0.72/0.72/0.73 0.72/0.72/0.72 0.72/0.72/0.72
O+FM+CA 0.93/0.92/0.92 0.77/0.77/0.78 0.77/0.76/0.77 0.77/0.76/0.77
@@
Figure 2: Pairwise model comparison via the Wilcoxon signed-rank test. Each cell compares two models: Blue (row model significantly better), Peach (worse), and Red (no significant difference).

### v2

@@ 3.2 Training set-up
3.2.1 Conventional augmentation
We applied a randomized transformation pipeline that included rotation (±15°), horizontal flipping, translation (±10%), scaling (0.8–1.2×), and Gaussian noise addition ($\sigma$ = 0.001–0.01).
3.2.2 Stable diffusion augmentation
During training, each image was paired with its corresponding prompt and provided to the Stable Diffusion model. We used a fine-tuning setup with mixed precision (fp16), a resolution of 128×128, exponential moving average, gradient accumulation with four steps, and a batch size of 1, optimizing for 50,000 steps at a learning rate of 1e-5. During inference, we evaluated three approaches: 1) image and prompt to image (SD1), 2) prompt to image (SD2), image to image (SD3) and 3) mask and prompt to image (SD4).
3.2.3 Flow matching augmentation
For FM, the model was optimized using the Adam optimizer with a learning rate of 1e-4 for 200 epochs. Our approach leverages FM with optimal transport to align predicted data flows with the target distribution. Class conditioning is achieved by incorporating a one-hot encoded vector via cross-attention, while mask conditioning is implemented through a parallel control network integrated via residual connections. During inference, we assessed guided generation using a combination of mask and class conditioning.
This resulted in 18 distinct training strategies for the ResNet18 classifier: 1) O, 2) CA, 3) SD1, 4) SD2, 5) SD3, 6) SD4, 7) FM, 8) O+CA, 9) O+SD1, 10) O+SD2, 11) O+SD3, 12) O+SD4, 12) O+FM, 14) O+CA+SD1, 15) O+CA+SD2, 16) O+CA+SD3, and 17) O+CA+SD4, 18) O+CA+FM.
3.2.4 ResNet18
We trained a ResNet18 classifier, initializing it with ImageNet1K pre-trained weights. The first convolutional layer was modified to retain a 3×3 kernel, and the fully connected layer was replaced with a dropout layer (0.2) followed by a linear layer matching the number of classes. The model was trained for 300 epochs using the Adam optimizer (learning rate = 1e-4, weight decay = 1e-5) with a cross-entropy loss function. A learning rate scheduler (ReduceLROnPlateau) was applied, reducing the rate by a factor of 0.8 if validation loss plateaued for 10 epochs, with a minimum learning rate of 2e-5. Furthermore, the training and validation split was set at a 9:1 ratio with stratification.
@@ 3.3.1 Augmentation metrics
Fréchet Inception Distance (FID) [22] and Kernel Inception Distance (KID) [23] metrics have been used for class-wise and overall comparisons between generated and original images. In both cases, 200 images per class were randomly selected from each dataset for evaluation.
3.3.2 Classification metrics
@@
The classification performance was evaluated on an external dataset using metrics, including precision, recall, F1-score, accuracy, and the area under the Receiver Operating Characteristic curve (ROC AUC). Given that it was a multiclass task, we applied various averaging techniques encompassing micro, macro, and weighted to provide a thorough evaluation across all classes.
3.3.3 Statistical method
@@
We compared different strategies using bootstrapping with 1,000 repetitions and sampling with replacement. Accuracy distributions were analyzed, and pairwise comparisons were made using the Wilcoxon rank sum test, considering p-values ¡ 0.05 as statistically significant.
3.3.4 GradCam
@@
We used Gradient-weighted Class Activation Mapping (GradCAM) on the trained ResNet18 model to generate heatmaps, visualizing the image regions that influenced the classifier’s decisions.
4 Results
@@
Figure 2 presents one sample per class from the original dataset alongside examples generated by different augmentation methods used in this study. Additionally, it includes samples from the external dataset with their corresponding Grad-CAM visualizations using the O+FM+CA model, highlighting the model’s focus during prediction.
Figure 2: Examples of original and augmented images for each class using different methods. Grad-CAM visualizations from the O+FM+CA model are also shown on external dataset samples, highlighting the model’s focus during prediction. (DG: Diffuse Goiter, NG: Nodular Goiter, NL: Normal, TI: Thyroiditis, O: Original, CA: Conventional Augmentation, SD: Stable Diffusion, FM: Flow Matching, EX: External)
Table 2 reports class-wise and different averaging of precision, recall, F1-score, and AUC for each model variant. Among all configurations, O+FM+CA achieves the best overall performance, with high F1-scores across all classes (0.77/0.75/0.60/0.94) and top micro/macro/weighted F1 (0.77/0.76/0.77) and strong AUC values (0.93/0.92/0.92). O+FM, without CA, also performs strongly (F1: 77/75/60/94; AUC: 95/93/94), slightly outperforming O+FM+CA in AUC but showing a more modest macro F1. Notably, adding CA slightly lowers AUC but improves balance across class performance. FM alone already provides strong gains over O alone, especially in the NL and TI classes, raising NL F1 from 0.39 to 0.51 and TI from 0.93 to 0.85.
Among the SD-based methods SD1 shows the strongest standalone performance among SD variants, with solid macro F1 (0.75) and consistent gains across all classes. When combined with O (O+SD1), it improves macro F1 from 0.69 (O) to 0.72. SD2 is competitive when combined with O, with O+SD2 achieving F1 of 0.72/0.71/0.54/0.93 and macro AUC of 0.92. However, its standalone version (SD2) performs poorly in NL and TI classes. SD3 underperforms across the board. Its class-wise F1 scores (0.51/0.29/0.27/0.59) and overall metrics (macro F1: 0.42; AUC: 0.67) suggest limited utility in isolation. SD4 achieves high precision in the TI class (0.58), but extremely low recall (0.35 in NG) leads to imbalanced performance. Despite strong DG precision (0.88), its macro F1 (0.46) and micro accuracy (0.59) are among the lowest.
Averaged metrics further reinforce the superiority of FM-based methods. O+FM yields the highest micro/macro/weighted precision, recall, and F1 (all 0.78 or 0.77), outperforming even the best SD combinations. SD1 and SD2, when combined with O and CA, reach macro F1 scores of 0.73–0.74 but still fall short of FM-based configurations.
Table 2: Classification model performance metrics by class and average metrics for each augmentation method ($\times 10^{-2}$).
Method
Precision
(DG/NG/NL/TI)
Recall
(DG/NG/NL/TI)
F1-score
(DG/NG/NL/TI)
AUC
(mic/mac/wei)
Precision
(mic/mac/wei)
Recall
(mic/mac/wei)
F1-score
(mic/mac/wei)
O 70/65/53/96 72/86/31/91 71/74/39/93 92/91/91 73/71/72 73/70/73 73/69/71
CA 75/70/55/92 62/77/62/97 68/73/58/95 93/92/92 74/73/75 74/74/74 74/74/74
SD1 80/74/53/92 72/71/59/98 76/72/56/95 93/92/92 76/74/76 76/75/76 76/75/76
SD2 76/57/45/95 70/68/53/74 73/62/49/83 88/87/88 67/68/70 67/66/67 67/67/68
SD3 55/27/32/54 48/32/23/66 51/29/27/59 68/67/67 44/42/44 44/42/44 44/42/43
SD4 88/66/42/58 44/35/64/100 58/46/50/74 82/86/86 59/63/67 59/61/59 59/57/58
FM 70/54/48/96 67/64/54/76 69/59/51/85 89/88/88 66/67/69 66/65/66 66/66/67
O+CA 73/59/61/99 72/83/41/89 72/69/49/94 92/92/92 73/73/74 73/71/73 73/71/73
O+SD1 74/67/55/96 70/85/46/89 72/75/50/92 92/92/92 74/73/74 74/73/74 74/72/74
O+SD2 78/60/61/98 67/87/49/89 72/71/54/93 92/92/92 74/74/76 74/73/74 74/73/74
O+SD3 72/62/57/93 71/83/39/87 72/71/46/90 90/90/90 72/71/72 72/70/72 72/70/71
O+SD4 76/62/59/95 66/87/47/89 71/73/53/92 92/92/92 73/73/74 73/72/73 73/72/73
O+FM 74/74/68/93 81/76/54/95 77/75/60/94 95/93/94 78/77/78 78/76/78 78/77/78
O+SD1+CA 75/64/54/96 66/82/50/91 70/72/52/93 93/92/93 73/72/74 73/72/73 73/72/73
O+SD2+CA 74/65/62/94 66/82/54/91 70/73/58/92 92/92/92 74/73/74 74/73/74 74/73/74
O+SD3+CA 77/63/60/91 65/78/58/90 71/70/59/90 92/91/91 73/73/74 73/73/73 73/72/73
O+SD4+CA 72/63/58/95 63/82/53/90 67/72/55/92 92/91/92 72/72/73 72/72/72 72/72/72
O+FM+CA 79/68/63/97 73/84/58/90 76/75/61/93 93/92/92 77/77/78 77/76/77 77/76/77
Figure 3 presents a pairwise comparison of different models using the Wilcoxon rank sum test, highlighting significant performance differences. The results demonstrate that O+FM and its extended variant (O+CA+FM) achieve consistently superior performance compared to most other models. Notably, SD3 and SD4 exhibit the weakest performance, being significantly outperformed by almost all other models. Additionally, O+SD1, O+SD2, and their CA-augmented counterparts exhibit stronger results than SD2, SD3, and SD4 alone, further validating the impact of structured generation methods, particularly when coupled with O and CA. This analysis reinforces the effectiveness of FM and structured data augmentation.
@@
Figure 3: Pairwise model comparison via the Wilcoxon signed-rank test. Each cell compares two models: Blue (row model significantly better), Peach (worse), and Red (no significant difference). (DG: Diffuse Goiter, NG: Nodular Goiter, NL: Normal, TI: Thyroiditis, O: Original, CA: Conventional Augmentation, SD: Stable Diffusion, FM: Flow Matching)
Table 3 highlights clear differences in generative quality across augmentation methods, based on FID and KID scores. FM achieves the best overall performance, with the lowest FID (0.66) and KID (0.83), indicating that its generated images closely align with the real data distribution across all classes. Among the SD-based methods, SD1 and SD2 show moderate performance. SD2 achieves a lower overall FID (3.88) than SD1 (4.17), but its KID is slightly higher (2.61 vs. 4.99), and its performance on the TI class is notably poor (KID of 30.02). SD1 is more stable across classes, with no extreme outliers, though its FID and KID are consistently higher than FM. SD3 outperforms SD4 overall, with a lower FID (2.77 vs. 17.99) and KID (4.66 vs. 33.59). However, SD3 still lags behind FM, particularly in the TI class (FID: 4.15 vs. FM’s 1.97). SD4 performs the worst by a significant margin—especially in the NG, NL, and TI classes—indicating high visual artifacts and weak alignment with real data. For example, its TI KID reaches 50.50, far exceeding all others. Across all classes, FM consistently produces the lowest FID and KID, with the best performance in NG (FID: 0.74) and DG (FID: 0.96), and strong results even in the more challenging NL and TI categories. In contrast, the SD methods vary widely, with no single SD variant performing best across all classes. This underscores FM’s robustness and the instability of diffusion-based augmentation without careful conditioning.
Table 3: Comparison of FID and KID scores across synthetic data generation methods.
Class SD1 SD2 SD3 SD4 FM
FID $\downarrow$ KID $\downarrow$ FID $\downarrow$ KID $\downarrow$ FID $\downarrow$ KID $\downarrow$ FID $\downarrow$ KID $\downarrow$ FID $\downarrow$ KID $\downarrow$
DG 4.75 6.77 2.94 2.31 1.56 1.94 10.30 15.40 0.96 0.80
NG 6.22 10.28 3.27 2.66 2.95 4.41 22.07 38.34 0.74 0.95
NL 4.56 7.14 1.75 2.05 3.39 5.68 18.68 36.74 0.85 2.08
TI 2.73 3.01 9.24 30.02 4.15 8.11 22.37 50.50 1.97 3.39
Overall 4.17 4.99 3.88 2.61 2.77 4.66 17.99 33.59 0.66 0.83
Standard deviations are less than 1e-5 for all measurements.
5 Discussion
Several studies have used diffusion-based models for data augmentation, leading to improved performance in classification tasks. However, only a few have incorporated physician reports as prompts to guide the diffusion process, which they reported to further enhance results. While all these works employed diffusion models, none explored the use of FM, which offers potential advantages in terms of efficiency and image quality.
Zhang et al. [8] investigated SinDDM [24], a single-image denoising diffusion model, to augment lung ultrasound data. They also introduced FewDDM, an extension trained on limited samples, which outperformed single-image GANs in generating high-quality synthetic images. Augmenting with SinDDM notably improved pathology classification, especially for minority classes. Despite generating less detailed images, FewDDM surpassed SinDDM and SinGAN in downstream performance by capturing local structural variations. The study highlighted that combining synthetic and CA techniques yielded the best classification results.
Hajianfar et al. [12] investigated the effectiveness of SD [17] as an advanced augmentation method in enhancing deep learning models for classifying scintigraphic thyroid images. They used reports from physicians without specific cleaning as prompts in the augmentation process. The SD, O, and CA were used to train a ResNet101V2 classifier in different scenarios. The results demonstrated that models trained with synthetic data achieved consistently better performance.
Balla et al. [25] explored strategies to address data scarcity in musculoskeletal US for osteoarthritis detection. They used CA with diffusion-based image synthesis, and the results showed that synthetic images generated through diffusion models retained anatomical fidelity and improved model generalization diagnostic accuracy, while CA sometimes hindered performance, highlighting the potential of using synthetic images.
Akrout et al. [13] advances data augmentation using text-to-image diffusion models to enhance a macroscopic skin disease dataset. By using text prompts, they gain fine-grained control over the image generation process. The results show that this generative augmentation approach maintains classification accuracy even when trained on a fully synthetic dataset.
FM [14], as a novel, more robust, and memory-efficient method for image synthesis has shown superiority over GANs and DDPMs. While it has not been widely used in the medical domain, it has demonstrated clear advantages.
In this study, we employed a variety of strategies for advanced augmentation. Specifically, we used image masks as conditions for both DDPM and FM, and we incorporated physicians’ reports as prompts for DDPM to maximize the available information for augmentation. Using FM, our objective was to improve both the efficiency and quality of image synthesis, making it a key component of our approach.
The comparison between O+FM and O+FM+CA reveals a key trade-off between overall accuracy and fairness across class distributions. While O+FM achieves the highest micro-averaged F1-score, O+FM+CA ensures more stable performance across all classes by improving macro and weighted scores. This suggests that if the goal is to optimize pure accuracy, O+FM should be preferred; however, if the objective is to achieve a more balanced model that is not overly biased toward some specific classes, O+FM+CA might be the better choice.
Among the SD-based models, SD1 which uses both the image and prompt during inference, consistently outperforms SD2, SD3, and SD4, indicating that combining visual and textual information leads to higher-quality synthetic data and improved classification performance. This advantage may stem from two factors: either the SD model inherently performs better when both image and prompt are available, especially the ones were existed in the training process, or the prompts themselves, derived from physician reports, carry clinically valuable context that enhances generation. In this study, the latter appears especially plausible, as these prompts are grounded in real diagnostic language specific to thyroid scintigraphy, potentially guiding the model to produce more realistic and relevant variations.
SD3, which performs image-to-image translation without textual input, ranks below SD1 and SD2. While it benefits from the visual structure of the original image, the lack of prompt input may limit its ability to generate semantically diverse or diagnostically meaningful augmentations. SD2, which performs prompt-only generation, achieves competitive overall FID but shows instability across classes, particularly with high KID in the TI category. This suggests that, while prompts can guide generation, relying on them alone may not provide enough structural information, especially for visually complex classes. SD4 performs the worst overall. It uses prompt and mask inputs during inference, but since masks were not present during training, the model lacks the capacity to meaningfully interpret them. As a result, the generated outputs are less coherent and lead to poor downstream performance.
Compared to the study by Hajianfar et al. [12], which evaluated multiple augmentation strategies including Stable Diffusion (similar to SD1 in our study), CA, and their combinations, our work introduces a more advanced augmentation pipeline by incorporating FM, a technique not examined in their framework, along with three additional SD-based model variants. Both studies support the benefit of combining synthetic and real data over using original images alone. However, our results demonstrate that FM consistently generates higher-fidelity images, as evidenced by superior FID and KID scores, which correlate with improved micro- and macro-averaged F1-scores. While they identified the SD1+O as the most effective approach, our FM-based methods, particularly O+FM and O+FM+CA—achieved better class-wise balance and generalization, supported by statistical analysis and Grad-CAM visualizations on external data. Furthermore, implementing SD3 allows us to evaluate the added value of prompts, which has not been done in their study.
The strong correlation between image quality metrics (FID and KID) and classification performance highlights the clinical importance of generating realistic synthetic data. FM consistently achieved the lowest FID and KID scores, reflecting its ability to produce high-fidelity images that enhance model generalization and reliability. In contrast, SD3 and SD4’s poor image quality was associated with degraded classification performance, illustrating that not all augmentation methods are beneficial. High-quality, distribution-aligned synthetic data is essential to avoid spurious correlations and ensure model trustworthiness. Additionally, Grad-CAM visualizations on external data confirm that models trained with FM focus on anatomically and pathologically relevant regions, supporting interpretability and clinical acceptance. These findings emphasize the viability of FM-based augmentation—particularly when combined with conventional techniques as a practical tool for improving AI-assisted diagnosis.
The data in this study is limited to a single ethnic group; to ensure clinical applicability, further validation on more diverse datasets is necessary. Incorporating US images commonly used in these patients alongside nuclear medicine images could potentially enhance classification performance. Moreover, the use of reports from both imaging modalities, namely US and nuclear medicine, to generate prompts for SD-based models may further improve the augmentation results.